OpenStudy (anonymous):

1/2(4x-8)=3x.. can anyone solve for x?

OpenStudy (anonymous):

x=-4??

OpenStudy (anonymous):

1/2(4x - 8) = 3x 2x - 4 = 3x -2x -2x x = -4